• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 16:01:42 on August12,2001

    Entry number 70671

    keyword=Day shift summary

    Shift summary

    8:00 data taking, runs 2825/21825
    10:30 MCC called, lost all RF, expected downtime 1-2 hours
    10:40 NMR P=35.4% (lasers off)
    10:42 problems ending run on left arm, restart CODA
    10:52 EPR P = 36.6% (3 long. lasers on)
    11:05 Tanja rebooted ROC2.
    11:07 Target in BeO position, moved l/2(t) plate out.
    11:55 started cosmics run
    12:25 MCC had a CHL trip -> no beam on this shift anymore.

    16:00 again: no beam!!!!

    Run summary
    Run #.................... good events
    2825/21825...............400k/386k
    2826/21826...............386k/370k
    2827/21827...............340k/335k
    2828/21828...............403k/392k
    2829/21829...............22k/47k (problems ending run on left arm)
    2830/21830...............cosmics